OpenStudy (anonymous):

Convert 77 degrees in Fahrenheit to Celsius using the formula F=9/5C+32. 72°C 61°C 35°C 25°C

OpenStudy (marigirl):

77=(9/5)c+32 first lets minus 32 from both side. can u write the equation now

OpenStudy (anonymous):

ya \

OpenStudy (marigirl):

now we are at 45=(9/5)c now divide both sides by (9/5)

OpenStudy (anonymous):

ok

OpenStudy (marigirl):

what is ur final answer

OpenStudy (anonymous):

25

OpenStudy (marigirl):

awesome.

OpenStudy (anonymous):

Which of the following equations is linear? y = |x| y = 2x x 2 + 5 = y y =1/3 x

OpenStudy (marigirl):

thanx

OpenStudy (marigirl):

y = |x| first one is an absolute value function and therefore not linear y=2x is a linear equation i hope the third one is x^2+5=y which will make it a quadratic, therefore, not linear y=(1/3)x is not linear.

OpenStudy (anonymous):

What are the next three terms in the sequence 11, 18, 25, 32, ...? 37, 44, 51 39, 46, 53 41, 48, 55 43, 50, 57

OpenStudy (anonymous):

@marigirl

OpenStudy (marigirl):

sry cant reply back now, but i can do it later on today. let me know if u still want the answer

OpenStudy (anonymous):

no I dont want answers I just want to know how to do it @marigirl